OpenStudy (anonymous):

check please find the distance between point P(-2,3) and Q (5,4), leave as simplified radical this is what i did P.Q=√(5+2)^2+(4-3)^2 P,Q=√49+1 P,Q=√50

OpenStudy (mathstudent55):

The distance formula is \[d = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\] for points \[(x_1, y_1) \] and \[(x_2, y_2)\]

OpenStudy (anonymous):

i did use the distance formula

OpenStudy (mathstudent55):

\[d = \sqrt{(5 + 2)^2 + (4 - 3)^2}\]

OpenStudy (mathstudent55):

\[d = \sqrt{49 + 1} = \sqrt{50}\]

OpenStudy (anonymous):

d=√50

OpenStudy (anonymous):

so what i did was right? i mean we cant solve the 50 further it would be decimals

OpenStudy (mathstudent55):

\[\sqrt{50} = \sqrt{25 \times 2} = \sqrt{25} \times \sqrt{2} \]

OpenStudy (anonymous):

√50=√(25 x 2)=5√2

OpenStudy (mathstudent55):

\[d = 5\sqrt{2}\]

OpenStudy (mathstudent55):

What you did is correct. You just didn't finish. You need to simplify the radical by taking the square root of the largest perfect square factor of 50 which is 25.

OpenStudy (anonymous):

ohhh so we keep simplifying lke 2 *25 and then we break that down i see how u guys did that now

OpenStudy (anonymous):

i know i thoought √50 would be the final answer i didnt know i could go further

OpenStudy (mathstudent55):

The problem asks for simplified radical. There is a step between sqrt(50) and a decimal. That is sqrt(50) = 5*sqrt(2)

OpenStudy (anonymous):

you can get the square root of 25 and leave the 2 inside radical because they are a product. And that's your simplified answer

OpenStudy (anonymous):

alright i get it thanks alot guys :)